# Modeling phase transformations in Mn-rich disordered rocksalt cathodes with charge-informed machine-learning interatomic potentials

### v1

@@ Abstract
Mn-rich disordered rocksalt (DRX) cathode materials exhibit a phase transformation from a disordered to a partially disordered spinel-like structure (-phase) during electrochemical cycling. In this computational study, we used charge-informed molecular dynamics with a fine-tuned CHGNet foundation potential to investigate the phase transformation in LixMn0.8Ti0.1O1.9F0.1. Our results indicate that transition metal migration occurs and reorders to form the spinel-like ordering in an FCC anion framework. The transformed structure contains a higher concentration of non-transition metal (0-TM) face-sharing channels, which are known to improve Li transport kinetics. Analysis of the Mn valence distribution suggests that the appearance of tetrahedral Mn2+ is a consequence of spinel-like ordering, rather than the trigger for cation migration as previously believed. Calculated equilibrium intercalation voltage profiles demonstrate that the -phase, unlike the ordered spinel, exhibits solid-solution signatures during the 0-TM to Li conversion reaction. A higher Li capacity is obtained than in the DRX phase. This study provides atomic insights into solid-state phase transformation and its relation to experimental electrochemistry, highlighting the potential of charge-informed machine learning interatomic potentials for understanding complex oxide materials.
@@ I Introduction
The increasing demand for sustainable energy storage has driven the development of high-energy-density and cost-effective cathode materials for rechargeable Li-ion batteries [1, 2, 3]. Disordered rocksalt (DRX) cathodes using manganese (Mn) as the earth-abundant redox-active cation have the potential to scale Li-ion energy storage to several TWh/year [4]. Recent studies demonstrate that initiating a Mn-rich disordered rocksalt structure undergoes an in-situ transformation into a partially disordered spinel-like phase (termed as -phase) upon electrochemical cycling, resulting in increased capacity and rate capability for cathode designs [5, 6, 7, 8, 9]. This approach mitigates the typical low capacity in other DRX cathodes by promoting a new voltage feature around 4 V and an extended 3 V plateau. Moreover, in contrast to conventional LixMn2O4 spinel cathodes, which suffer from the two-phase reactions leading to inhomogeneity-induced stresses and mechanical degradation between , the -phase maintains a single (solid-solution) phase reaction throughout the cycling process. Cai et al. [9] demonstrates that such a series of -phase material exhibits an impressive capacity of over 280 mAh/g and an energy density exceeding 800 Wh/kg while maintaining minimal voltage decay even after extensive cycling.
@@
The spinel is distinct from the disordered rocksalt based on cation ordering in an FCC anion framework. Figure 1a–1c presents the cation orderings in a conceptualized graph representation as a projection of the 3-dimensional FCC structure shown in Fig. 1c, including the disordered rocksalt (Fmm), spinel (Fdm), and partially disordered spinel (-phase) structures. The green and purple spheres represent Li and Mn ions, respectively, occupying the octahedral sites (, denoted as ) of the FCC rocksalt framework. For simplicity, the stoichiometric LiMnO2 is used for illustration in Fig. 1a-c. The graph edges connect the sites to the interstitial tetrahedral sites (, denoted as ), which are represented by white dashed circles in Fig. 1. In the spinel LiMnO2 ordering, the octahedral sites are categorized into and sites as shown in Fig. 1b. All sites are occupied by Li, while all sites are occupied by Mn. The tetrahedral site neighbored by four Li is denoted as an site in a spinel. When the spinel is delithiated to Li0.5MnO2, the sites become vacancies and the sites are all occupied by Li. More generally, the tetrahedral site neighbored by four Li is also termed a 0-TM channel. Urban et al. [10] demonstrated that long-range connected 0-TM percolating channels are essential for facile Li diffusion in an FCC framework, with the spinel ordering having the highest amount of 0-TM channels in theory. The partially disordered spinel is conceptually illustrated in Fig. 1c as an intermediate state between the spinel and disordered ordering.
@@
The phase transformation from DRX to -phase is related to the transition metal (TM) migration of the mobile Mn ion when neighboring cation vacancies are created [11]. The TM can take the pathway connected by the octahedral-tetrahedral-octahedral (--) sites in an FCC framework. For the migration process, it is typically concluded that Mn migration is strongly coupled with its charge state, Mn4+ is generally considered immobile, while Mn3+ and Mn2+ are considered more mobile according to the nudged-elastic band calculations [12, 13, 14]. However, simulating these charge-coupled dynamics has long been challenging, as the required time and length scales are inaccessible to conventional ab initio methods. Previously, cluster expansion (CE) methods have been shown successfully as a surrogate model of density functional theory (DFT) calculations for modeling the configurational thermodynamics in complex oxides via Monte Carlo (MC) simulation at scale [15, 16, 17]. Nonetheless, the limited expressibility and coarse-grained nature of these models render them less adequate for capturing the complex charge-coupled dynamical processes involved in TM migrations [18, 9]. The advent of machine learning interatomic potentials (MLIPs) has opened new avenues for addressing these problems by enabling large-scale and nano-second-long simulations with ab-initio quantum accuracy [19, 20, 21]. By using magnetic moments as a proxy for atomic charge, Deng et al. [21] recently simulated a phase transformation from an orthorhombic to a spinel-like phase in Li0.5MnO2 with charge-informed MD. It identified the timescale discrepancies between charge disproportionation and ion hops, suggesting that the emergence of Mn2+ is correlated to long-range spinel-like ordering but not solely to the local hopping into the tetrahedral occupancy.
@@
In this study, we extend the computational study of the phase transformation in LixMn0.8Ti0.1O1.9F0.1 (LMTOF), which is a Mn-rich DRX cathode. We parametrized a CHGNet-MLIP using r2SCAN-DFT calculations for various atomic configurations to sample the potential energy surface in this chemical space. The fine-tuned CHGNet-MLIP was used to implement charge-informed MD simulations starting from a disordered phase in Li0.6Mn0.8Ti0.1O1.9F0.1. The MD simulations revealed a phase transformation from a disordered rocksalt to a spinel-like () phase. From the magnetic moment predictions, we demonstrate that the Mn migration is mainly contributed by Mn3+ at the beginning, while Mn2+ collectively emerges as the spinel-like ordering forms. Furthermore, we investigated the impact of this partial disorder on the electrochemical intercalation voltage profiles based on the structures obtained from the MD simulations. Our work provides insights into the application of MLIPs for studying complex energy materials and offers a mechanistic understanding of the partial cation-disorder effect on intercalation electrochemistry at the atomic scale.
@@ II.1 Machine learning interatomic potential
The nodes are regularized by DFT magnetic moments to carry rich information about both local ionic environments and to serve as a proxy for the atomic charge. The forces and stress are calculated via auto-differentiation of the total energy with respect to the atomic Cartesian coordinates and strain:
This approach allows CHGNet to capture the complex relationships between charge distributions and energetics with given atomic configurations.
@@ II.2 Potential energy surface sampling
While CHGNet [21] as a pre-trained foundation potential achieves reasonable performance in predicting materials stability on out-of-distribution datasets [23], recent benchmarks observed a systematic softening of the ML predicted potential energy surface (PES) compared to the DFT ground truth, particularly for atomic configurations from MD simulations with high energy [24, 25]. To improve simulation accuracy, we performed multi-faceted sampling of the PES for the Li–Mn–Ti–O–F chemical space [26], as follows:
@@
Exisiting dataset – We collected the compounds in the Materials Project database within the Li–Mn–Ti–O–F chemical space that have decomposition energy (energy above the convex hull, ) lower than 0.1 eV/atom as a general coverage of such chemical space [27]. Furthermore, we adopted the structures from the training dataset of previously reported cluster expansion (CE) models for DRXs containing Mn2+/Mn3+ as redox-active species (DRX-CE) [28, 29]. All these structures in the DRX-CE dataset have rocksalt lattices but different cation and anion orderings. These structures were relaxed by DFT calculations. The relaxed structures with rocksalt frameworks were used to fit a cluster expansion Hamiltonian (lattice model fitting, see Appendix B). The relaxation trajectories from structures from the two datasets were used to fine-tune a Trial CHGNet.
@@
Partially disordered structures – Given that existing datasets do not include atomic configurations specifically representing partially disordered structures with spinel-like orderings, we developed a custom sampling procedure to generate these configurations. (1) Monte Carlo (MC) simulations were employed to obtain the disordered rocksalt atomic configurations equilibrate at finite temperatures, which were generated in supercells of the primitive spinel cell (see Appendix B for details).
Given these disordered structures, we compared the occupancy of the spinel supercell and the disordered rocksalt in such a supercell lattice. The octahedral cation sites are categorized into and sites. Under this classification, a random distribution of TM yields the disordered structure, while an increased number of TM occupations on sites indicates more spinel-like ordering.
@@
To generate different degrees of spinel order, we applied site exchanges between the two nearest octahedral sites to transform the disordered structure towards the spinel ordering through either (a) an exchange between a TM and a Li via a Li3M tetrahedron; or (b) an exchange between a TM and a Li. The (a) step is to create more spinel-like ordering and the (b) step is to exchange the cation ordering within the sites to potentially create more available Li3M tetrahedron. We assume that TM migration occurs through the -- path with no other TM neighbors in the delithiated states (i.e., through the Li3M tetrahedron, Fig. 3). By iteratively applying site exchanges within different Li3M tetrahedra, one can obtain a collection of structures ranging from 0% (disordered) to nearly 100% spinel ordering.
@@
TM migration pathways – Since TM migration is a rare, high-energy-barrier event [9], especially for unfavorable valence states like Mn4+, we implemented a passive sampling approach to effectively model the diffusion kinetics with the following steps:
@@
For each site exchange step via the Li3M tetrahedron, three different states are generated as A (initial state, TM occupied on site), B (intermediate state, TM occupied on site), and C (final state, TM occupied on site), which are shown in Fig. 3. The other neighboring Li atoms are removed to create vacancies, for which we assume no face-sharing between the migrating TM and cations. The structures in the A/B/C states were relaxed using the Trial-CHGNet fitted with structures from the existing datasets. After structure relaxation, five images interpolated along A-B (- path) and B-C (- path) are evaluated with static DFT calculations, analogous to the intermediate states for nudged elastic band (NEB) calculations. Furthermore, we follow the same approach to generate the direct migration from the - path (A-C) for labeling high-energy and unfavorable states.
@@
In addition, we considered two different delithiation levels for sampling TM migration pathways: (a) A dilute level, where only three Li ions are removed to create a diffusion channel for TM migration (e.g., a pristine supercell structure with Li44Mn32Ti4O76F4 yields a delithiated structure with Li41Mn32Ti4O76F4). This level allows for the investigation of local TM hopping energetics. (b) A medium delithiation level, where 0.5 Li per formula unit (f.u.) is removed from the structure to create delithiated states (e.g., Li ions are randomly removed to create a structure with Li24Mn32Ti4O76F4). This level allows the sampling of delithiated states, representing the charging process in a battery. By sampling partially disordered structures at these two delithiation levels, we include a more comprehensive description of the cation ordering energetics, TM hop, and relevant local environments in which TM hopping occurs at the composition that is energetically favorable for spinel formation (i.e., the composition close to Li0.5TMO2).
By adding the static DFT calculations of these sampled atomic configurations to the relaxation trajectories, another MLIP was parameterized (denoted as Sample-CHGNet) to perform MD simulations for enhanced sampling.
@@
MD trajectories – To further sample the thermally excited atomic configurations at finite temperatures, we employed Sample-CHGNet to perform molecular dynamics (MD) simulations with disordered structures across various compositions. These structures were first generated in the pristine (lithiated) state from the CE-MC simulations at K using supercells of and formula units (f.u.), including Li1.2Mn0.5Ti0.3O1.9F0.1, Li1.2Mn0.6Ti0.2O1.8F0.2, Li1.1Mn0.8Ti0.1O1.9F0.1, and Li1.1Mn0.7Ti0.2O2.0. For each disordered structure, three different delithiated levels (removing 0.2/0.4/0.6 Li per f.u. from the pristine structure) were generated to initiate the MD simulations using the Sample-CHGNet. The MD simulations were run at K for ns with a time step of fs under NVT ensembles. From each MD trajectory, 200 structures were randomly sampled, and their energies were calculated using static DFT and added to the training set. In addition to the training set, we further ran CHGNet-MD simulations in NVT ensembles at K using Sample-CHGNet with a supercell of Li44-xMn32Ti4O76F4 ( f.u.). These trajectories were further evaluated by r2SCAN-DFT static calculations and were set as the test set.

### v2

# Modeling phase transformations in Mn-rich disordered rocksalt cathodes with machine learning interatomic potentials

@@ Abstract
Mn-rich disordered rocksalt (DRX) cathode materials exhibit a phase transformation from a disordered to a partially disordered spinel-like structure (-phase) during electrochemical cycling. In this computational study, we used charge-informed molecular dynamics with a fine-tuned CHGNet foundation potential to investigate the phase transformation in LixMn0.8Ti0.1O1.9F0.1. Our results indicate that transition metal migration occurs and reorders to form the spinel-like ordering in an FCC anion framework. The transformed structure contains a higher concentration of non-transition metal (0-TM) face-sharing channels, which are known to improve Li transport kinetics. Analysis of the Mn valence distribution suggests that the appearance of tetrahedral Mn2+ is a consequence of spinel-like ordering, rather than the trigger for cation migration as previously suggested. Calculated equilibrium intercalation voltage profiles demonstrate that the -phase, unlike the ordered spinel, exhibits solid-solution signatures at low voltage. A higher Li capacity is obtained than in the DRX phase. This study provides atomic insights into solid-state phase transformation and its relation to experimental electrochemistry, highlighting the potential of machine learning interatomic potentials for understanding complex oxide materials.
@@ I Introduction
The increasing demand for sustainable energy storage has driven the development of high-energy-density and cost-effective cathode materials for rechargeable Li-ion batteries [1, 2, 3]. Disordered rocksalt (DRX) cathodes using manganese (Mn) as the earth-abundant redox-active cation have the potential to scale Li-ion energy storage to several TWh/year [4]. Recent studies demonstrate that Mn-rich disordered rocksalt structures undergo an in-situ transformation into a partially disordered spinel-like phase (termed the -phase) upon electrochemical cycling [5, 6, 7, 8, 9], which results in increased capacity and rate capability for cathode materials. This approach mitigates the typical low capacity in other DRX cathodes by promoting a new voltage feature around 4 V and an extended 3 V plateau. Moreover, in contrast to conventional LixMn2O4 spinel cathodes, which suffer from a two-phase reaction when fully lithiated between , leading to inhomogeneity-induced stresses and mechanical degradation, the -phase maintains a single (solid-solution) phase reaction throughout the cycling process. Cai et al. [9] demonstrates that such -phase material exhibits an impressive capacity of over 280 mAh/g and an energy density exceeding 800 Wh/kg while maintaining minimal voltage decay even after extensive cycling.
@@
The spinel is distinguished from the disordered rocksalt by its cation ordering in the common FCC anion framework. Figure 1a–1c presents the cation orderings in a conceptualized graph representation as a projection of the 3-dimensional FCC structure shown in Fig. 1d, including the disordered rocksalt (Fmm), spinel (Fdm), and partially disordered spinel (-phase) structures. The green and purple spheres represent Li and Mn ions, respectively, occupying the octahedral sites (, denoted as ) of the FCC rocksalt framework. For simplicity, the stoichiometric LiMnO2 is used for illustration in Fig. 1a-c. The graph edges connect the sites to the interstitial tetrahedral sites (, denoted as ), which are represented by white dashed circles in Fig. 1. In the spinel LiMnO2 ordering, the octahedral sites are categorized into and sites as shown in Fig. 1b. All sites are occupied by Li, while all sites are occupied by Mn. Note that we refer to spinel LiMnO2 here as the product of lithiation of the traditional spinel LiMn2O4. The tetrahedral site neighbored by four Li is denoted as an site in a spinel. When the spinel is delithiated to Li0.5MnO2, the sites become vacant and the sites are occupied by Li, representing a traditional spinel structure. More generally, the tetrahedral site neighbored by four Li is also termed a 0-TM channel. Urban et al. [10] demonstrated that long-range connected 0-TM percolating channels are essential for facile Li diffusion in an FCC framework, with the spinel ordering having the highest amount of 0-TM channels of any ordering. The partially disordered spinel is conceptually illustrated in Fig. 1c as an intermediate state between the spinel and the disordered rocksalt.
@@
The phase transformation from the DRX to -phase is related to the transition metal (TM) migration of the mobile Mn ion when neighboring cation vacancies are created [11]. The TM migration can occur along the pathway connected by the octahedral-tetrahedral-octahedral (--) sites in an FCC framework. The Mn migration barrier is strongly coupled with its charge state: Mn4+ is generally considered immobile, while Mn3+ and Mn2+ are considered more mobile according to the nudged-elastic band calculations [12, 13, 14]. However, simulating these charge-coupled dynamics has long been challenging, as the required time and length scales are inaccessible to conventional ab initio methods. Previously, cluster expansion (CE) methods have successfully acted as a surrogate model to density functional theory (DFT) calculations for modeling the configurational thermodynamics in complex oxides via Monte Carlo (MC) simulation at scale [15, 16, 17]. Nonetheless, the limited expressibility and coarse-grained nature of these models render them less adequate for capturing the complex charge-coupled dynamical processes involved in TM migrations [18, 9]. The advent of machine learning interatomic potentials (MLIPs) has opened new avenues for addressing these problems by enabling large-scale and nano-second-long simulations with ab-initio quantum accuracy [19, 20, 21]. By using magnetic moments as a proxy for atomic charge, Deng et al. [21] recently simulated a phase transformation from an orthorhombic to a spinel-like phase in Li0.5MnO2 with charge-informed MD. It identified the timescale discrepancies between charge disproportionation and ion hops, suggesting that the emergence of Mn2+ is correlated to long-range spinel-like ordering but not solely to the local hopping into the tetrahedral occupancy.
@@
In this study, we study the complex phase transformation in LixMn0.8Ti0.1O1.9F0.1 (LMTOF), which is a Mn-rich DRX cathode material. We parametrized a CHGNet-MLIP using r2SCAN-DFT calculations for various atomic configurations to sample the potential energy surface in this chemical space. The fine-tuned CHGNet-MLIP was used to implement charge-informed MD simulations starting from a disordered phase in Li0.6Mn0.8Ti0.1O1.9F0.1. The MD simulations reveal a phase transformation from a disordered rocksalt to a spinel-like () phase. From the magnetic moment predictions, we demonstrate that the Mn migration mainly occurs as Mn3+ in the initial part of the transformation, while Mn2+ collectively emerges as the spinel-like ordering forms. Furthermore, we investigated the impact of this partial disorder on the electrochemical intercalation voltage profiles using the structures obtained from the MD simulations. Our work provides insights into the application of MLIPs for studying complex energy materials and offers a mechanistic understanding of the partial cation-disorder effect on intercalation electrochemistry at the atomic scale.
@@ II.1 Machine learning interatomic potential
The nodes are regularized by DFT magnetic moments to carry rich information about both local ionic environments. The valence state of TM is inferred from the predicted magnetic moment as a proxy of the atomic charge, which is a direct output of the MLIP rather than an input parameter. The forces and stress are calculated via auto-differentiation of the total energy with respect to the atomic Cartesian coordinates and strain:
This approach allows CHGNet to capture the complex relationships between charge distributions and energetics with given atomic configurations, which is essential to properly model the kinetic path of evolving TM oxides [23].
@@ II.2 Potential energy surface sampling
CHGNet [21] as a pre-trained foundation potential achieves reasonable performance in predicting materials stability on out-of-distribution datasets [24]. However, recent benchmarks observed a systematic softening of the ML predicted potential energy surface (PES) compared to the DFT ground truth, particularly for atomic configurations with high energy [25, 26]. To improve simulation accuracy, we performed multi-faceted sampling of the PES for the Li–Mn–Ti–O–F chemical space [27], as follows:
@@
Exisiting dataset – We collected the compounds in the Materials Project database within the Li–Mn–Ti–O–F chemical space that have decomposition energy (energy above the convex hull, ) lower than 0.1 eV/atom as a general coverage of such chemical space [28]. Furthermore, we adopted the structures from the training dataset of previously reported cluster expansion (CE) models for disordered rocksalts containing Mn2+/Mn3+ as redox-active species (DRX-CE) [29, 30]. All these structures in the DRX-CE dataset have rocksalt lattices but different cation and anion orderings. These structures were relaxed by DFT calculations. The relaxed structures with rocksalt frameworks were used to fit a cluster expansion Hamiltonian (lattice model fitting, see Appendix B). The relaxation trajectories from structures from the two datasets were used to fine-tune a Trial CHGNet.
@@
Partially disordered structures – Given that existing datasets do not include atomic configurations specifically representing partially disordered structures with spinel-like orderings, we developed a custom sampling procedure to generate these configurations. (a) Monte Carlo (MC) simulations were employed to obtain disordered rocksalt atomic configurations in a supercell of the primitive spinel cell by equilibration at high temperatures (see Appendix B). (b) The octahedral cation sites are then categorized into and sites by overlaying the template of the perfectly ordered spinel. Under this classification, a random distribution of TM yields the disordered structure, while preferential TM occupation of sites indicates more spinel-like ordering.
@@
To generate different degrees of spinel order, we applied site exchanges between two nearest neighbor octahedral sites to transform the disordered structure towards the spinel ordering through: (a) An exchange between a TM and a Li in a Li3M tetrahedron. We selected a Li3M tetrahedron that contains three Li atoms (including at least one Li) and one metal atom. A site exchange is applied to the Li and the TM. (b) An exchange between a TM and a Li. Step (a) increases the spinel-like order, but it may halt if no suitable Li3M tetrahedra are available to facilitate the exchange. To address this, Step (b) rearranges cations within the sites. While this internal exchange does not change the overall degree of spinel ordering, its purpose is to create new Li3M tetrahedra, thereby enabling further transformations via Step (a). By iteratively applying site exchanges within different Li3M tetrahedra, one can obtain a collection of structures ranging from disordered to spinel ordering.
@@
TM migration pathways – TM migration is a rare, high-energy-barrier event [9], especially for unfavorable valence states such as Mn4+. To effectively model the diffusion kinetics, we implemented a passive sampling approach with the following steps:
@@
A TM can only migrate from its octahedral site into or through a tetrahedral site which has no other face-sharing TMs. In the fully lithiated state, there are Li3M tetrahedra where the tetrahedron is surrounded by only one TM and 3 Li. For each migration event via a Li3M tetrahedron, three different states are generated as A (initial state, TM occupied on site), B (intermediate state, TM occupied on site), and C (final state, TM occupied on site), which are shown in Fig. 3. The Li atoms face-sharing with the site are removed to create vacancies so that there are no face-sharing ions with the migrating TM and cations. The structures in the A/B/C states were relaxed using the Trial-CHGNet. After structure relaxation, five images interpolated along A-B (- path) and B-C (- path) are evaluated with static DFT calculations, analogous to the intermediate states for nudged elastic band (NEB) calculations. Furthermore, we follow the same approach to generate states along a direct migration between sites, which cuts through their shared edge (A-C) so that high-energy states are considered in the MLIP fitting [31].
@@
In addition, we considered two different delithiation levels for sampling TM migration pathways: (a) A dilute level, where only three Li ions are removed to create a diffusion channel for TM migration (e.g., a pristine supercell structure with Li44Mn32Ti4O76F4 yields a delithiated structure with Li41Mn32Ti4O76F4). This level allows for the investigation of the environment dependence of the TM hopping energetics. (b) A medium delithiation level, where 0.5 Li per formula unit (f.u.) is removed from the structure to create delithiated states (e.g., Li ions are randomly removed to create a structure with Li24Mn32Ti4O76F4). This level allows the sampling of delithiated states, representing the charging process in a battery. By sampling partially disordered structures at these two delithiation levels, we include a more comprehensive description of the cation ordering energetics, TM hop, and relevant local environments in which TM hopping occurs at the composition that is energetically favorable for spinel formation (i.e., the composition close to Li0.5TMO2) [32]. By adding the static DFT calculations of these sampled atomic configurations to the relaxation trajectories, another MLIP was parameterized (denoted as Sample-CHGNet) to perform MD simulations for enhanced sampling.
@@
MD trajectories – To further sample the thermally excited atomic configurations at finite temperatures, we employed Sample-CHGNet to perform molecular dynamics (MD) simulations on disordered structures across various compositions. These structures were first generated in the pristine (lithiated) state from the CE-MC simulations at K using supercells of and formula units (f.u.), including Li1.2Mn0.5Ti0.3O1.9F0.1, Li1.2Mn0.6Ti0.2O1.8F0.2, Li1.1Mn0.8Ti0.1O1.9F0.1, and Li1.1Mn0.7Ti0.2O2.0. For each disordered structure, three different delithiated levels (removing 0.2/0.4/0.6 Li per f.u. from the pristine structure) were generated to initiate the MD simulations using the Sample-CHGNet. The MD simulations were run at K for ns with a time step of fs under NVT ensembles. From each MD trajectory, 200 structures were randomly sampled, and their energies were calculated using static DFT and added to the training set. In addition to the training set, we further ran CHGNet-MD simulations in NVT ensembles at K using Sample-CHGNet with a supercell of Li44-xMn32Ti4O76F4 ( f.u.). These trajectories were further evaluated by r2SCAN-DFT static calculations and were set as the test set.